C-5 Sheet Metal Trades- The length of a rectangular piece of sheet metal must be 5 in. longer than the width. Find the exact dimension that will provide an area of 374 square inches.
Answer by scott8148(6628)

let x=width, so x+5=length ___ area=L*W ___ 374=x(x+5) ___ 374=x^2+5x

x^2+5x-374=0 ___ (x+22)(x-17)=0